Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
Six years ago
Bob = 4Mary - 6
In four years
Bob = 2Mary + 4
4M - 6 = 2M + 4
2M = 10
Mary = 5 years old
Substitute both equations with 5, and Bob is 14
I just want to know if that is right or wrong?

Found 2 solutions by oscargut, stanbon:
Answer by oscargut(2103)

Six years ago bob was four times as old as Mary. In four years he will be twice as old as Mary. How old are they?
B= Bob's age now
M= Mary's age now
You did it wrong you have to start like this:
B-6=4(M-6)
B+4=2(M+4)
solving we have
B=26 and M=11

Answer by stanbon(75887)

Six years ago bob was four times as old as Mary.
In four years he will be twice as old as Mary.
How old are they?
-----------
Let Mary's age now be "m".
Let Bob's age now be "b".
---------------------------
6 yrs. ago equation:
(b-6) = 4(m-6)
------------------------------
4 yrs. from now equation:
b = 2m
--------------------------
Substitute to solve for "m".
(2m-6) = 4(m-6)
2m-6 = 4m-24
2m = 18
m = 9 yrs old (Mary's age now)
-------------------
Since b = 2m, b = 2*9 = 18 yrs old (Bob's age now)
===========================
Cheers,
Stan H.